Please help me solve this problem! thank you!
solve each quadratic equation :

16v^2 = 49

or
15s^2 - 8s - 12 = 0

either problem would be a great help! thanks

Answer by lwsshak3(11628)

16v^2 = 49
Take sqrt of both sides
4v=±√49=±7
v=7/4 or-7/4
..
15s^2 - 8s - 12 = 0
factor
(5s-6)(3s+2)=0
5s-6=0
s=6/5
or
3s+2=0
s=-2/3